OpenStudy (anonymous):

Quadrilateral QRST is a rectangle. RT = 5x + 1 and QS = 6x – 3. Find QS

OpenStudy (strawberry17):

that is not what I got...

OpenStudy (anonymous):

Now I got 4 is that what you got?

OpenStudy (strawberry17):

yes! that is right answer :)

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

Quadrilateral WXZY is a square. If YX = 30 and WT = 2x – 7, then find the value of x. @strawberry17

OpenStudy (anonymous):

I got 18.5

OpenStudy (strawberry17):

since WT is only half of YX, then divide YX in half and solve: 2x - 7 = 15

OpenStudy (anonymous):

So eleven?

OpenStudy (strawberry17):

yes :)

OpenStudy (crissy15):

Quadrilateral QRST is a rectangle. RT = 5x + 1 and QS = 6x – 3. Find QS. RT = QS 5x + 1 = 6x – 3 4 = x QS = 6(4) – 3 = 21